An element with mass 630 grams decays by

18.8% per minute. How much of the element

is remaining after 7 minutes, to the nearest

10th of a gram?


1
Expert's answer
2022-04-16T04:13:01-0400

"m=m_0(1-0.188)^7=630\cdot (1-0.188)^7=241.5\space g."